# Strategies for Pretraining Neural Operators

### v1

@@ Appendix C Comparison of Downstream Dataset Size
Figure 3: Fixed Future Scaling Behavior: For each model, a specific PDE/distribution is displayed. Within each graph, the performance of various pretraining strategies at different sample sizes is displayed. Validation errors are averaged over 5 seeds, and error bars denote 1 standard deviation. Derivative errors are omitted as outliers.
@@
Figure 4: Auto-regressive Scaling Behavior: For each model, a specific PDE/distribution is displayed. Within each graph, the performance of various pretraining strategies at different sample sizes is displayed. Validation errors are averaged over 5 seeds, and error bars denote 1 standard deviation.
@@ D.2 Model Details
We implement modern FNO and Unet architectures according to Gupta & Brandstetter (2022). Furthermore, we implement DeepONet architectures according to DeepXDE (Lu et al., 2021b), and use the original implementation for OFormer (Li et al., 2023a). The hyperparameters used for the models are described in Table 9(d).
@@
Figure 5: t-SNE Embeddings of CV Pretrained Models: We display latent embeddings after pretraining models on Binary, TimeSort. or Jigsaw objectives. We see that models learn to sort/classify sorted samples well and can visualize the relative difficulties of the proposed pretraining strategies.
@@ D.3 Pretraining Details
During pretraining, different strategies require different implementations and hyperparameters. A consideration is that many models need a linear head during pretraining to project model outputs to the classification or regression dimension. Since models will be used for physics prediction, their outputs will be in the shape of the solution field, rather than cross entropy probabilities or regressed values. We use a lightweight CNN projector to downsample and flatten model outputs to the desired dimension. Models are generally trained for 200 epochs with a batch size of 32 using Adam, weight decay, and a OneCycle scheduler for five seeds.
@@
Jigsaw is implemented similarly to other sorting frameworks, however due to sorting along multiple axes the number of possible shuffled sequences quickly increases. We mitigate this by using spatial and temporal patches to ensure a sequence length of 8 when shuffled, resulting in $8!=40320$ possible permutations. This is still a large number of classes for a task, therefore we deterministically choose 1000 samples with the largest Hamming distance between the shuffled sequence and original sequence. Contrary to the binary case, shuffled samples with larger Hamming distances are more challenging due to needing to sort more patches The CNN projection head is modified accordingly to output 1000 logits for a cross-entropy loss. In general, jigsaw sorting tends to be more challenging, however, models can still display reasonable performance; we display t-SNE embeddings of a pretrained FNO model on a validation set of jigsaw shuffled samples in Figure 5(c).
@@ D.5 Fine-tuning Details
During fine-tuning, models trained until convergence for fixed-future or auto-regressive prediction and repeated for five seeds. In fixed-future prediction, models are given the solution field at $t=[0,8)$ and the target is at $t=32$. For auto-regressive prediction, models are given the solution field at $t=[0,8)$ and the target is at $t=[8,16)$. After this prediction, the models use their own output to predict the next step $t=[16,24)$ until the time horizon of $t=32$. To stabilize auto-regressive rollout, we implement temporal bundling and the pushforward trick (Brandstetter et al., 2023). Losses are calculated using a relative L2 norm (Li et al., 2021); validation losses are averaged across batch size and accumulated over timesteps or, in the case of fixed-future prediction, at only one timestep. For experiments with different fine-tuning sample sizes, samples are randomly chosen from 1024 possible samples to reach the desired number of samples for each seed. We use an Adam optimizer with weight decay and a CosineAnnealing scheduler. All experiments are run on a NVIDIA GeForce RTX 2080Ti GPU.

### v2

@@ 5.4 Generalization Behavior
6 Discussion
6.1 Theoretical Insights
There are many insights from the broader deep learning community that can help inform our results. In our study, an important factor in pretraining performance is the choice of model architecture, which is reflected both in both PDE and deep learning literature. Many recent PDE works have chosen to include attention modules in their architectures (Hao et al., 2024; Herde et al., 2024; Li et al., 2023a; McCabe et al., 2023; Hang et al., 2024), which is supported by the observation that transformer architectures can be expressive by making the least assumptions on the structure of data (Jiang et al., 2022). We similarly find that OFormer tends to display larger improvements compared to other architectures. Furthermore, Kolesnikov et al. (2020) find that training on larger datasets is vital for transferability, which we also observe by limiting the fine-tuning data.
Additionally, the use of data augmentations can be understood as assembling a more comprehensive training set such that the distance between the training set and validation/test sets is reduced (Shorten & Khoshgoftaar, 2019). In the context of PDEs, physics data can contain different coefficients, initial conditions, and boundary conditions; augmentations can help to mimic invariances to these conditions (e.g., spatially shifting data to model different initial conditions, scaling data to mimic different coefficients, etc.) such that models can perform well in spite of these challenges. The utility of data augmentation is further exacerbated by the difficulty in generating high-quality PDE data from numerical simulation, which could motivate future work in exploring more complex data augmentation approaches such as generating synthetic samples from GANs (Sandfort et al., 2019) or neural style transfer (Gatys et al., 2016).
Lastly, we consider the connection of different PDE pretraining strategies to broader deep learning pretraining insights. In general, it is understood that unsupervised learning is beneficial due to the abundance of unlabeled data and the human labor required to label samples (Bengio et al., 2021). However, in the context of PDE data, the goal is usually to predict a future frame of a current solution or to upsample low-resolution physics, both of which do not require human intervention to label data. In addition, numerical simulation is required to generate PDE data, which is the limiting factor in assembling large datasets. As a result, we hypothesize that this limitation is the main factor as to why unsupervised PDE methods relying on surrogate objectives do not perform as well as supervised pretraining, also known as transfer learning. Currently, unsupervised methods rely on the same data as supervised methods, both because of the computational cost of numerical simulation and because of the lack of human labor needed to label supervised samples. Addressing this limitation could be a promising direction for future work on unsupervised PDE learning.
6.2 Best Practices
Our experiments have demonstrated that transfer learning combined with the shift augmentation performs best in the majority of our experiments. For pretraining strategies, we believe that transfer learning performs best because additional pretraining strategies are not able to extract additional, task-specific information better than transfer learning. Having identical training tasks is simply more effective in this case. For data augmentation, the shift augmentation is the only augmentation that preserves the underlying dynamics of our data since all of our systems are spatially invariant. We believe that this improves performance over the other augmentation strategies, which do not preserve the underlying dynamics. For further applications, transfer learning + shift augmentation is a good initial pretraining strategy, but practitioners will likely need to tune and adapt additional strategies to achieve maximal performance. Furthermore, previous work (Brandstetter et al., 2022) has demonstrated improved performance by combining multiple augmentations, which may help across many systems. However, due to computational cost, we leave this to future work for each practitioner’s specific application, as best practices are likely model and dataset dependent. Lastly, to maximize the benefit of pretraining, we recommend collecting as much pretraining data as possible.
@@ Appendix C Comparison of Downstream Dataset Size
Figure 3: Fixed Future Scaling Behavior: For each model, a specific PDE/distribution is displayed. Within each graph, the error of various pretraining strategies at different sample sizes is displayed. Validation errors are averaged over 5 seeds, and error bars denote 1 std-dev. Derivative errors are omitted as outliers.
@@
Figure 4: Auto-regressive Scaling Behavior: For each model, a specific PDE/distribution is displayed. Within each graph, the performance of various pretraining strategies at different sample sizes is displayed. Validation errors are averaged over 5 seeds, and error bars denote 1 std-dev.
@@ D.2 Model Details
Figure 5: t-SNE Embeddings of CV Pretrained Models: We display latent embeddings after pretraining models on Binary, TimeSort. or Jigsaw objectives. We see that models learn to sort/classify shuffled samples well and can visualize the relative difficulties of the proposed pretraining strategies.
We implement modern FNO and Unet architectures according to Gupta & Brandstetter (2022). Furthermore, we implement DeepONet architectures according to DeepXDE (Lu et al., 2021b), and use the original implementation for OFormer (Li et al., 2023a). The hyperparameters used for the models are described in Table 9(d). For FNO, we use the standard spectral convolution and residual layers, as well as additional input channels for the x, y, and time coordinates and multiple timesteps. The Unet implementation uses multiple ConvNext blocks followed by an upsample/downsample block at each layer. Additionally, an attention block is used at the middle layer between upsampling and downsampling. When implementing OFormer, an attention module is used to aggregate information from the inputs as well as output query coordinates; this is aggregated into a latent encoding which is propagated with an MLP and decoded at future timesteps. Lastly, DeepONet uses an implementation that bundles multiple timesteps into both the branch and trunk net MLPs, these predictions are combined to ouput multiple future timesteps.
@@ D.3 Pretraining Details
During pretraining, different strategies require different implementations and hyperparameters. A consideration is that many models need a linear head during pretraining to project model outputs to the classification or regression dimension. Since models will be used for physics prediction, their outputs will be in the shape of the solution field, rather than cross entropy probabilities or regressed values. We use a lightweight CNN projector to downsample and flatten model outputs to the desired dimension. Models are generally trained for 200 epochs with a batch size of 32 using Adam with a learning rate of 1e-3, a weight decay of 1e-6, and a OneCycle scheduler for five seeds. A note is that for binary pretraining, the task is much easier and therefore the pretraining can be done with 100 epochs.
@@
Jigsaw is implemented similarly to other sorting frameworks, however due to sorting along multiple axes the number of possible shuffled sequences quickly increases. We mitigate this by using spatial and temporal patches to ensure a sequence length of 8 when shuffled, resulting in $8!=40320$ possible permutations. This is still a large number of classes for a task, therefore we deterministically choose 1000 samples with the largest Hamming distance between the shuffled sequence and original sequence. Contrary to the binary case, shuffled samples with larger Hamming distances are more challenging due to needing to sort more patches. The CNN projection head is modified accordingly to output 1000 logits for a cross-entropy loss. In general, jigsaw sorting tends to be more challenging, however, models can still display reasonable performance; we display t-SNE embeddings of a pretrained FNO model on a validation set of jigsaw shuffled samples in Figure 5(c).
@@ D.5 Fine-tuning Details
During fine-tuning, models trained until convergence for fixed-future or auto-regressive prediction and repeated for five seeds. In fixed-future prediction, models are given the solution field at $t=[0,8)$ and the target is at $t=32$. For auto-regressive prediction, models are given the solution field at $t=[0,8)$ and the target is at $t=[8,16)$. After this prediction, the models use their own output to predict the next step $t=[16,24)$ until the time horizon of $t=32$. To stabilize auto-regressive rollout, we implement temporal bundling and the pushforward trick (Brandstetter et al., 2023). Losses are calculated using a relative L2 norm (Li et al., 2021); validation losses are averaged across batch size and accumulated over timesteps or, in the case of fixed-future prediction, at only one timestep. For experiments with different fine-tuning sample sizes, samples are randomly chosen from 1024 possible samples to reach the desired number of samples for each seed. These fine-tuning datasets are independently resampled from the same pool for each seed, and larger datasets are not required to be strict supersets of the smaller datasets, which could increase the variance of results. We use an Adam optimizer with a learning rate of 1e-3, with weight decay of 1e-6, and a CosineAnnealing scheduler. All experiments are run on a NVIDIA GeForce RTX 2080Ti GPU.
Appendix E Auxiliary Results
During our experimentation, there were different pretraining strategies that did not improve model performance over a baseline of no pretraining. These are omitted from consideration in the results and analysis, however, for completeness, preliminary results that informed this decision are reported here.
E.1 Sorting Spatially Shuffled Sequences
In this section, we compare SpaceSort with a baseline without pretraining. Both models are fine-tuned on 500 samples that are within the pretraining distribution from the respective PDEs on a fixed-future prediction objective. We hypothesize that spatially shuffling PDEs may confuse models during pretraining since it destroys important information, as physical systems need to be continuous in space.
Table 10: Models are pretrained on 9216 combined 2D Heat, Advection, and Burgers samples and finetuned on 500 samples for each PDE. L2 errors (×$\times 10^{-3}$) are calculated on 256 validation samples and averaged over five seeds.
Table 11: Fixed-Future Pretraining Results with SpaceSort
PDE Model None SpaceSort
Heat FNO 0.75 2.61
DeepONet 2.13 2.39
OFormer 2.38 10.78
Unet 0.46 1.51
Adv FNO 11.04 19.17
DeepONet 30.95 31.17
OFormer 30.13 31.26
Unet 12.37 20.34
Burgers FNO 2.20 3.83
DeepONet 14.04 14.63
OFormer 5.99 15.70
Unet 3.21 4.53
E.2 Lie Contrastive Self-Supervised Learning
In this section, we compare VICRegBardes et al. (2022) against baseline FNO training for a fixed-future prediction task. We use values of $\nu=1$, $\mu=25$, $\lambda=25$, matching values from the original paper. In this case, we use the first 8 frames to predict the system state at frame 32.
Table 12: Models are pretrained on 9216 combined 2D Heat, Advection, and Burgers samples and finetuned on 100 samples from the Heat equation. L2 errores (×$\times 10^{3}$) are calculated on 256 validation samples and averaged over five seeds.
Table 13: Fixed-Future Pretraining Results with VICReg
Dataset FNO FNO VICReg
Heat 1.45 8.03